# A Watermark for Black-Box Language Models

### v1

@@ Acknowledgements
The authors would like to thank Rob McAdam, Abhi Shelat, Matteo Frigo, Ashkan Soleymani, John Kirchenbauer, and Mukund Sundararajan for insightful discussions without which this work would not be possible.
@@ A.2 Omitted Experimental Results
Figure 2 shows the effect of varying the amount of random token corruption on detection pAUC. We observe the same trend as for AUC. Figure 3 plots a histogram of the entropy of the underlying next-token probability distribution under temperature 1 random sampling without watermarking across our dataset. We see the entropy is concentrated between 0.5 and 3 nats. We plot the AUC lower bound predicted by Theorem 4.2 ($k=1,m=1024$) sweeping our entropy term $\alpha$ across this range, with the understanding that for sufficiently large $m$, our $\alpha$ is a good estimator of the true underlying entropy. In Figure 4 we look at the impact of $m$ and $T$ on our AUC bound when the optimal $\alpha=\log(m)$ is plugged in. We see sharp diminishing returns w.r.t. $m$ (performance saturates after around $m=10$ for all $T$’s). We empirically observe this saturation in Table 1, where AUC saturates at 97.7% at $m=16$ — that is, increasing $m$ beyond 16 has negligible impact. Furthermore, we observe that the bound increases sharply with $T$, corroborating the trend we see empirically in Figure 1.
@@
Table 6: Main results (mixed $T$’s for AUC and pAUC where max FPR is 1%) for Gemma-7B-instruct on the eli5-category test split. AUC and pAUC are scaled by 100. We observe the same trends here as with Mistral-7B-instruct on databricks-dolly-15k. When $k=1$ and $m=1024$ (white-box setting) we are slightly better in perplexity and detection (sans corruption) than Kuditipudi et al. (2023) and on-par with Aaronson (2023). Kirchenbauer et al. (2023a) can always outperform on detection by cranking up $\delta$, but when matched on perplexity, we achieve better detection. For example, $\delta=0.5$ gives perplexity of 1.649 and AUC of 61.6% whereas we achieve perplexities / AUC’s of 1.610 and 93.2% when $k=1,m=1024$ and even 1.645 / 89.7% when $k=50,m=16$ (black-box).
@@
Table 7: Detection performance (mixed $T$’s) when a sum-based $p$-value is used in the length correction of Aaronson (2023). We observe slightly worse performance than using Fisher’s method to combine the $p$-values of individual tests. AUCs and pAUCs are scaled by 100. For paraphrasing, target lengths of {150, 200, 250} are used in computing AUC and pAUC.
@@ A.3 Omitted Proofs
Now, we prove that the outputs are identically distributed with the same distribution as their inputs. To do this, consider the $v$-th call in isolation and for ease of notation, let $\left\{Q_{1},\dots,Q_{m}\right\}=\mathbf{Q}_{v}$ and $X_{w}=\psi_{F,n,K,P}(\mathbf{Q}_{v})$. Let $\{(X_{1},c_{1}),\dots,(X_{j},c_{j})\}$ be the unique sequences and corresponding counts. Note that the $\{(X_{i},c_{i})\}_{i}$ need not be independent (it is easy to come up with a counter-example). Let $S_{i}$ be the integer seeds for $X_{i}$ after deduplication. Conditioned on $(c_{1},\dots,c_{j})$, $\{S_{i,l}\}_{i,l}$ is independent and so $\{R_{i,l}\}_{i,l}$ consists of i.i.d. draws from $F$ by virtue of pseudorandomness. As $F$ is also continuous, we have that when conditioned on $(c_{1},\dots,c_{j})$, $u_{i}\overset{iid}{\sim}U(0,1)$ for $i=1,\dots,j$, by the inverse-sampling theorem.
Let $x$ be any sequence. We wish to show that ${\mathbb{P}}(X_{w}=x)=\mu(x)$. Let $c=\sum_{i}\mathbf{1}[Q_{i}=x]$. The independence of the $\mathbf{1}[Q_{i}=x]$’s follows from the independence of the $Q_{i}$’s, and thus $c\sim\text{Binomial}(m,\mu(x))$. Clearly, (${\mathbb{P}}(\{x\;\text{selected}\}\;|\;c=0)=0$. If $c>0$ then obviously one of the $X_{i}$’s is $x$, and we can, without loss of generality, label $X_{1}=x$ and $c_{1}=c$, so that (${\mathbb{P}}(\{x\;\text{selected}\}\;|\;c=i)={\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i)$. Now,
(${\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1},\dots,c_{j})$ =($={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}u_{t}^{m/c_{t}}\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
=($={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}\frac{\log(u_{t})}{c_{t}/m}\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
=($={\mathbb{P}}\left(\left\{1=\operatorname{argmin}_{t}\log(-\log(u_{t}))-\log(c_{t}/m)\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
=($={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}-\log(-\log(u_{t}))+\log(c_{t}/m)\right\}\;\middle|\;c_{1},\dots,c_{j}\right).$
@@
(${\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i)$ $=\sum_{c_{2},\dots,c_{j}}\frac{{\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i,c_{2},\dots,c_{j}){\mathbb{P}}(c_{1}=i,c_{2},\dots,c_{j})}{{\mathbb{P}}(c_{1}=i)}$
@@
Suppose that regardless of the conditioning prompt, the LM outputs one of two sequences — $x_{1}$ or $x_{2}$ with equal probability. Let $u_{i}=\textsc{ScoreSeqs}(F,(x_{i}),K,n,P)$ for $i\in\{1,2\}$. If $m$ is very large, then it becomes very likely that $X_{1}=x_{1}$, $X_{2}=x_{2}$ (modulo the labeling) and $c_{1}\approx c_{2}\approx m/2$ and so $\operatorname{argmax}_{i=1}^{2}u_{i}^{m/c_{i}}\approx\operatorname{argmax}_{i}u_{i}$. The outputs to two sequential calls to WatermarkSingle should not be independent, because the output and key are dependent and the key is shared across calls. Concretely, if the output to the first call is $x_{1}$ we learn that our scheme with key $K$ prefers $x_{1}$ over $x_{2}$, and so we will likely output $x_{1}$ in the second call. In contrast, if we had not observed the first call (and our prior on the key had not been updated), we may have returned each sequence with equal probability.
@@
Distortion-free for WatermarkRecursive follows easily from induction on $t$, the number of keys (and hence the number of recursive calls). When $t=1$, the LM is the true underlying language model, so the outputs are i.i.d. from $\mu$. We get $t=v+1$ by combining Lemma A.1 with the inductive step — that the outputs of WatermarkRecursive with keys $(K_{2},\dots,K_{v+1})$ are i.i.d. from $\mu$.
@@
To do this, we introduce two sets of random variables: $\{X_{u}^{(i)}\}_{i=1}^{\infty}$ represents $k$-sized chunks of the model’s response when watermarking is not employed — that is, $X_{u}^{(i)}$ represents non-watermarked response tokens for indices $(i-1)k+1$ to $ik$. Unused chunks can be set to a sentinel value like $\phi$. $\{X_{w}^{(i)}\}_{i}$ represents the same collection but when Watermark is employed. Let $x$ be a sequence of any length. Partition $x$ into contiguous $k$-sized chunks $(x_{1},\dots,x_{t})$. Note that $x_{t}$ may have length less than $k$ if the stop-token was reached in that chunk, but all other chunks have exactly $k$ tokens. With $P$ as the original prompt, we need to show ${\mathbb{P}}(X_{w}=x\mid P)={\mathbb{P}}(X_{u}=x\mid P)$, where $X_{w}$ and $X_{u}$ are the watermarked and non-watermarked responses of any length.
@@
Differentiating this w.r.t to $t$, we recover the pdf of $\text{Beta}(m/c_{i^{*}},1)$. ∎
@@
$F$ is ${U(0,1)}$. The detection score is $F_{T}\left(\sum_{j}R_{j}\right)$ with $R_{j}\overset{iid}{\sim}F$ under $\mathcal{H}_{0}$ and when conditioned on the counts $C$ and the argmax token ids $I^{*}$, $R_{j}\sim\text{Beta}\left(m/C_{j,I^{*}_{j}},1\right)$ under $\mathcal{H}_{1}$. Redefine $s_{0}$ and $s_{1}$ to be $\sum_{j}R_{j}$ under $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ respectively.
@@
where $t\sim\text{IrwinHall}(T)$ since $s_{0}$ is the sum of $T$ i.i.d. ${U(0,1)}$’s. Our task now is to find a lower-bound for $s_{1}$. Noting independence across tokens and that $R_{j}\in[0,1]$, we can use Popoviciu’s bound on variance to obtain,
@@
Let $r$ be the PRF value for some $n$-gram from the text we wish to text. Let $F_{0}=-\text{Gamma}(1/k,\beta)$ with pdf $f_{0}$ and $F_{1}=-\text{Gamma}(1/k,m\beta)$ with pdf $f_{1}$. By definition, $r\sim F_{0}$ under $\mathcal{H}_{0}$. By our assumptions, $c_{i}=1$ and $|R_{i}|=k,\;\forall i$. So, $\operatorname{argmax}_{i=1}^{m}u_{i}^{m/c_{i}}=\operatorname{argmax}_{i=1}^{m}u_{i}=\operatorname{argmax}_{i}F_{k}\left(\sum_{j}R_{i,j}\right)=\operatorname{argmax}_{i}\sum_{j}R_{i,j}=\operatorname{argmin}_{i}-\sum_{j}R_{i,j}$, where the second-to-last equality follows from the monotonicity of $F_{k}$. $-\sum_{j}R_{i,j}\sim\text{Gamma}(k/k,\beta)=\text{Exp}(1,\beta)$. $\sum_{j}R_{i^{*},j}\sim-\text{Exp}(1,m\beta)$, because the minimum of Exponentials is Exponential. Thus, $\forall j,\;R_{i^{*},j}\sim-\text{Gamma}(1/k,m\beta)=F_{1}$ and $r\sim F_{1}$ under $\mathcal{H}_{1}$. Now let $R$ refer to the $T$ test-time PRF values. From the independence of test $n$-grams, the log-likelihood ratio test has score $s(R)=\sum_{i=1}^{T}\left(\log f_{1}(R_{i})-\log f_{0}(R_{i})\right)$ and the fact that it is the uniformly most powerful test follows directly from the Neyman–Pearson lemma. We now have that,

### v3

@@ 6 Conclusion
Broader Impacts
Our watermarking scheme, like other ones, can be used to understand the provenance of text — this can be both good and bad. On one hand, it can be used to catch inappropriate uses of a language model (e.g. illicit applications that violate terms of service), or help users understand whether they can trust a piece of text (if it is LLM-generated and its factuality is suspicious, it may be a hallucination). On the other hand, it may violate user privacy — LLM providers can see where users post the outputs of their models, and to make matters worse, users may not be able to detect if this is happening.
@@ Acknowledgements
The authors would firstly like to thank Rob McAdam, Abhi Shelat, and Matteo Frigo for their early, unpublished work on the topic and Sumanth Dathathri for critical technical feedback on it as well as on our work. They proposed and experimentally validated a white-box scheme whose encoding procedure looks identical to that of our flat one when $k=1$. Secondly, the authors would like to thank Ashkan Soleymani, John Kirchenbauer, and Mukund Sundararajan for insightful discussions that were instrumental in the development of this work.
@@ A.2 Omitted Experimental Results
Figure 2 shows the effect of varying the amount of random token corruption on detection pAUC. We observe the same trend as for AUC. Figure 3 plots a histogram of the entropy of the underlying next-token probability distribution under temperature 1 random sampling without watermarking across our dataset. We see the entropy is concentrated between 0.5 and 3 nats. We plot the AUC lower bound predicted by Theorem 4.2 ($k=1,m=1024$) sweeping our entropy term $\alpha$ across this range, with the understanding that for sufficiently large $m$, our $\alpha$ is a good estimator of the true underlying entropy. In Figure 4 we look at the impact of $m$ and $T$ on our AUC bound when the optimal $\alpha=\log(m)$ is plugged in. We see sharp diminishing returns with respect to $m$ (performance saturates after around $m=10$ for all $T$’s). We empirically observe this saturation in Table 1, where AUC saturates at 97.7% at $m=16$ — that is, increasing $m$ beyond 16 has negligible impact. Furthermore, we observe that the bound increases sharply with $T$, corroborating the trend we see empirically in Figure 1.
@@
Testing the independence assumption.
Central to our theory was the assumption that the seeds, deduplicated across candidate sequences, are independent. We now empirically test that assumption as follows. We watermark 20 random prompts from databricks-dolly-15k using Mistral-7B-instruct with a minimum and maximum text length of 250 and 300 tokens respectively. The flat scheme with $F=U(0,1)$ for various $k$, $m$, and most importantly $n$, is used. During the watermarking procedure we track the $u_{i}$ for each candidate sequence. We pool all these $u_{i}$’s together across decoding steps and prompts, and run a hypothesis test. Our null hypothesis is that all the seeds are independent. Under the null, the set of $u_{i}$’s should look like one giant i.i.d. sample from $U(0,1)$, and so we can apply a two-sided Kolmogorov-Smirnov (KS) test to observe how well our sample’s empirical CDF agrees with $U(0,1)$. Table 8 shows the statistic and $p$-value for various hyperparameter settings, while Figure 8 depicts the empirical CDFs themselves for visual confirmation. We clearly observe the importance of $n$ on the independence assumption. As we expect, decreasing $n$ means a greater the chance of reusing seeds and as a result implies less independence and more distortion. Recall, however, that the benefit of smaller $n$ is greater robustness to attacks, so there are trade-offs. The choice of $k$ has less of an effect than $n$.
@@
Table 6: Main results (mixed $T$’s for AUC and pAUC where max FPR is 1%) for Gemma-7B-instruct on the eli5-category test split. AUC and pAUC are scaled by 100. We observe the same trends here as with Mistral-7B-instruct on databricks-dolly-15k. When $k=1$ and $m=1024$ (white-box setting) we are slightly better in perplexity and detection (sans corruption) than kuditipudi2023robust and on-par with aaronson. kirchenbauer2023watermark can always outperform on detection by cranking up $\delta$, but when matched on perplexity, we achieve better detection. For example, $\delta=0.5$ gives perplexity of 1.649 and AUC of 61.6% whereas we achieve perplexities / AUC’s of 1.610 and 93.2% when $k=1,m=1024$ and even 1.645 / 89.7% when $k=50,m=16$ (black-box).
@@
Figure 8: Empirical CDF of the $u_{i}$’s tracked during watermarking 20 random prompts from databricks-dolly-15k using Mistral-7B-instruct with our flat scheme. We can clearly see that decreasing $n$ causes the empirical distribution to deviate further from the CDF of $U(0,1)$ (the line $y=x$).
@@
Table 7: Detection performance (mixed $T$’s) when a sum-based $p$-value is used in the length correction of aaronson. We observe slightly worse performance than using Fisher’s method to combine the $p$-values of individual tests. AUCs and pAUCs are scaled by 100. For paraphrasing, target lengths of {150, 200, 250} are used in computing AUC and pAUC.
$n$ $k$ $m$ # samples KS $p$-value
4 1 32 38605 0.0122 2.0e-05
3 1 32 38997 0.0124 1.3e-05
2 1 32 27055 0.0503 5.9e-60
1 1 32 21185 0.0920 2.4e-156
4 10 32 15655 0.0104 6.7e-02
Table 8: KS statistic and corresponding $p$-value for our test of independence. We see that decreasing $n$ results in a larger KS (more distortion) and a lower $p$-value for rejecting the null that the seeds are independent.
@@ A.3 Omitted Proofs
Now, we prove that the outputs are identically distributed with the same distribution as their inputs. To do this, consider the $v$-th call in isolation and for ease of notation, let $\left\{Q_{1},\dots,Q_{m}\right\}=\mathbf{Q}_{v}$ and $X_{w}=\psi_{F,m,K,P}(\mathbf{Q}_{v})$. Let $\{(X_{1},c_{1}),\dots,(X_{j},c_{j})\}$ be the unique sequences and corresponding counts. Note that the $\{(X_{i},c_{i})\}_{i}$ need not be independent (it is easy to come up with a counter-example). Let $S_{i}$ be the integer seeds for $X_{i}$ after deduplication. Conditioned on $(c_{1},\dots,c_{j})$, $\{S_{i,l}\}_{i,l}$ is independent and so $\{R_{i,l}\}_{i,l}$ consists of i.i.d. draws from $F$ by virtue of pseudorandomness. As $F$ is also continuous, we have that when conditioned on $(c_{1},\dots,c_{j})$, $u_{i}\overset{iid}{\sim}U(0,1)$ for $i=1,\dots,j$, by the inverse-sampling theorem.
Let $x$ be any sequence. We wish to show that ${\mathbb{P}}(X_{w}=x)=\mu(x)$. Let $c=\sum_{i}\mathbf{1}[Q_{i}=x]$. The independence of the $\mathbf{1}[Q_{i}=x]$’s follows from the independence of the $Q_{i}$’s, and thus $c\sim\text{Binomial}(m,\mu(x))$. Clearly, ${\mathbb{P}}(\{x\;\text{selected}\}\;|\;c=0)=0$. If $c>0$ then obviously one of the $X_{i}$’s is $x$, and we can, without loss of generality, label $X_{1}=x$ and $c_{1}=c$, so that ${\mathbb{P}}(\{x\;\text{selected}\}\;|\;c=i)={\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i)$. Now,
${\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1},\dots,c_{j})$ $={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}u_{t}^{m/c_{t}}\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
$={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}\frac{\log(u_{t})}{c_{t}/m}\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
$={\mathbb{P}}\left(\left\{1=\operatorname{argmin}_{t}\log(-\log(u_{t}))-\log(c_{t}/m)\right\}\;\middle|\;c_{1},\dots,c_{j}\right)$
$={\mathbb{P}}\left(\left\{1=\operatorname{argmax}_{t}-\log(-\log(u_{t}))+\log(c_{t}/m)\right\}\;\middle|\;c_{1},\dots,c_{j}\right).$
@@
${\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i)$ $=\sum_{c_{2},\dots,c_{j}}\frac{{\mathbb{P}}(\{X_{1}\;\text{selected}\}\;|\;c_{1}=i,c_{2},\dots,c_{j}){\mathbb{P}}(c_{1}=i,c_{2},\dots,c_{j})}{{\mathbb{P}}(c_{1}=i)}$
@@
Suppose that regardless of the conditioning prompt, the LLM outputs one of two sequences — $x_{1}$ or $x_{2}$ with equal probability. Let $u_{i}=\textsc{ScoreSeqs}(F,(x_{i}),K,n,P)$ for $i\in\{1,2\}$. If $m$ is very large, then it becomes very likely that $X_{1}=x_{1}$, $X_{2}=x_{2}$ (modulo the labeling) and $c_{1}\approx c_{2}\approx m/2$ and so $\operatorname{argmax}_{i=1}^{2}u_{i}^{m/c_{i}}\approx\operatorname{argmax}_{i}u_{i}$. The outputs to two sequential calls to WatermarkSingle should not be independent, because the output and key are dependent and the key is shared across calls. Concretely, if the output to the first call is $x_{1}$ we learn that our scheme with key $K$ prefers $x_{1}$ over $x_{2}$, and so we will likely output $x_{1}$ in the second call. In contrast, if we had not observed the first call (and our prior on the key had not been updated), we may have returned each sequence with equal probability.
@@
Distortion-free for WatermarkRecursive follows easily from induction on $t$, the number of keys (and hence the number of recursive calls). When $t=1$, the LLM is the true underlying language model, so the outputs are i.i.d. from $\mu$. We get $t=v+1$ by combining Lemma A.1 with the inductive step — that the outputs of WatermarkRecursive with keys $(K_{2},\dots,K_{v+1})$ are i.i.d. from $\mu$.
@@
To do this, we introduce two sets of random variables: $\{X_{u}^{(i)}\}_{i=1}^{\infty}$ represents $k$-sized chunks of the model’s response when watermarking is not employed — that is, $X_{u}^{(i)}$ represents non-watermarked response tokens for indices (i1)k+1(i-1)k+1 to $ik$. Unused chunks can be set to a sentinel value like $\phi$. $\{X_{w}^{(i)}\}_{i}$ represents the same collection but when Watermark is employed. Let $x$ be a sequence of any length. Partition $x$ into contiguous $k$-sized chunks $(x_{1},\dots,x_{t})$. Note that $x_{t}$ may have length less than $k$ if the stop-token was reached in that chunk, but all other chunks have exactly $k$ tokens. With $P$ as the original prompt, we need to show ${\mathbb{P}}(X_{w}=x\mid P)={\mathbb{P}}(X_{u}=x\mid P)$, where $X_{w}$ and $X_{u}$ are the watermarked and non-watermarked responses of any length.
@@
Differentiating this with respect to $t$, we recover the pdf of $\text{Beta}(m/c_{i^{*}},1)$. ∎
@@
$F$ is U(0,1){U(0,1)}. The detection score is $F_{T}\left(\sum_{j}R_{j}\right)$ with $R_{j}\overset{iid}{\sim}F$ under $\mathcal{H}_{0}$ and when conditioned on the counts $C$ and the argmax token ids $I^{*}$, $R_{j}\sim\text{Beta}\left(m/C_{j,I^{*}_{j}},1\right)$ under $\mathcal{H}_{1}$. Redefine $s_{0}$ and $s_{1}$ to be $\sum_{j}R_{j}$ under $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ respectively.
@@
where $t\sim\text{IrwinHall}(T)$ since $s_{0}$ is the sum of $T$ i.i.d. U(0,1){U(0,1)}’s. Our task now is to find a lower-bound for $s_{1}$. Noting independence across tokens and that $R_{j}\in[0,1]$, we can use Popoviciu’s bound on variance to obtain
@@
Let $r$ be the PRF value for some $n$-gram from the text we wish to text. Let $F_{0}=-\text{Gamma}(1/k,\beta)$ with pdf $f_{0}$ and $F_{1}=-\text{Gamma}(1/k,m\beta)$ with pdf $f_{1}$. By definition, $r\sim F_{0}$ under $\mathcal{H}_{0}$. By our assumptions, $c_{i}=1$ and $|R_{i}|=k,\;\forall i$. So, $\operatorname{argmax}_{i=1}^{m}u_{i}^{m/c_{i}}=\operatorname{argmax}_{i=1}^{m}u_{i}=\operatorname{argmax}_{i}F_{k}\left(\sum_{j}R_{i,j}\right)=\operatorname{argmax}_{i}\sum_{j}R_{i,j}=\operatorname{argmin}_{i}-\sum_{j}R_{i,j}$, where the second-to-last equality follows from the monotonicity of $F_{k}$. $\sum_{j}R_{i,j}\sim\text{Gamma}(k/k,\beta)=\text{Exp}(1,\beta)$. $\sum_{j}R_{i^{*},j}\sim-\text{Exp}(1,m\beta)$, because the minimum of Exponentials is Exponential. Thus, $\forall j,\;R_{i^{*},j}\sim-\text{Gamma}(1/k,m\beta)=F_{1}$ and $r\sim F_{1}$ under $\mathcal{H}_{1}$. Now let $R$ refer to the $T$ test-time PRF values. From the independence of test $n$-grams, the log-likelihood ratio test has score $s(R)=\sum_{i=1}^{T}\left(\log f_{1}(R_{i})-\log f_{0}(R_{i})\right)$ and the fact that it is the uniformly most powerful test follows directly from the Neyman–Pearson lemma. We now have that,